OpenStudy (anonymous):

consider a closed rectangular box with a square base, whose edges have length x. If x is measured with error at most 2% and the height y is measured with error at most 3%, use a differential to estimate the corresponding %error in computing the box's surface area.

OpenStudy (anonymous):

how much of an estimate can we make? I used x+dx, y+dy and found a term for the error, but it doesnt exactly come out nice.

OpenStudy (anonymous):

the answer should be +/- 5%

OpenStudy (anonymous):

ok I got about 4%

OpenStudy (anonymous):

here is my attempt, given: dx/x=2%, dy/y=3%, and i am looking for ds/s

OpenStudy (anonymous):

using partial derivatives, ds=(4x+4y)dx+(4x)dy

OpenStudy (anonymous):

ds/s=(4x+4y)dx+(4x)dy /(2x^2+4xy), and then i am stuck.. i was able to get dx/x but not dy/y

OpenStudy (anonymous):

k give me a sec to think about this

OpenStudy (anonymous):

Thanks so much!!!!

OpenStudy (anonymous):

k it turns out the same way as the way I did it actually: multiply the first part of the numerator by x/x and the second part by y/y then you have ((4x^2+4xy)dx/x+4xy(dy/y)) we know dx/x=.02, dy/y=.03 pluging in (.08x^2+.08xy+.12xy)/(2x^2+4xy) =(.08x^2+.2xy)/(2x^2+4xy) now here is why I asked how big of an estimate we can make, we round .08 to .1 so that it comes out nicely: .1(x^2+2xy)/(2(x^2+2xy))=.1/2=.05 5%

OpenStudy (anonymous):

AWESOME! i got stuck because i didn't round it off.......

OpenStudy (anonymous):

yeah it's not nice, but they just want an estimate so I geuss it is ok

OpenStudy (anonymous):

Thank you very much :)

OpenStudy (anonymous):

you're welcome :)